OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

y-4=3/4(x+8) y-4=3/4x+6 y=3/4x+10

OpenStudy (anonymous):

Thank you!! :-) @WordGEEK